OpenStudy (anonymous):

Multiply and write the result in standard form: (10 - 8i)(2 - 3i)

OpenStudy (anonymous):

\[(10-8i)(2-3i)=(10*2)+(-8i*2)+(-3i*10)+(-8i*-3i)=20-16i-30i+24i^2\] \[20-16i-30i+24i^2=20-46i+24i^2\] \[i^2=1 \rightarrow 20-46i+24i^2=20-46i+24(-1)=20-46i-24=-4-46i\] (-4-46i) is the standard form.

OpenStudy (anonymous):

any question?